### v3

# RiteWeight: Randomized Iterative Trajectory Reweighting for Steady-State Distributions Without Discretization Error

Abstract
A significant challenge in molecular dynamics (MD) simulations is ensuring that sampled configurations converge to the equilibrium or nonequilibrium stationary distribution of interest. Lack of convergence constrains the estimation of free energies as well as rates and mechanisms of molecular transitions. Here, we introduce the “Randomized ITErative trajectory reWeighting” (RiteWeight) algorithm to estimate a stationary distribution from unconverged simulation data. This method iteratively reweights trajectory segments in a self-consistent way by solving for the stationary distribution of a Markov state model (MSM), updating segment weights, and employing a new random clustering in each iteration. The repeated clustering mitigates the phase-space discretization error inherent in existing trajectory reweighting techniques and yields quasi-continuous configuration-space distributions. RiteWeight accurately recovers the stationary distribution, even without requiring the Markov property at the cluster level. We present mathematical analysis of the RiteWeight fixed point. We empirically validate the method using both synthetic MD Trp-cage trajectories, for which the stationary solution is exactly calculable, and standard atomistic MD Trp-cage trajectories extracted from a long reference simulation. In both test systems, RiteWeight corrects flawed distributions and generates accurate observables for equilibrium and nonequilibrium steady states. The results highlight the value of correcting the underlying trajectory distribution rather than using a standard MSM.
Significance
Molecular dynamics (MD) simulation is a key tool for studying the behavior of proteins and other biomolecules, but despite four decades of hardware and algorithm advances, MD cannot completely characterize the biomolecular behavior of most systems of interest. Typically, the molecular configurations generated by MD and related methods fail to conform to the equilibrium or nonequilibrium steady-state distribution of interest, thereby limiting the accuracy of computed observables such as rate constants. The present report introduces a novel approach for correcting mis-distributed configurations using a Randomized ITErative reWeighting (RiteWeight) strategy. The approach is validated for protein folding systems under both equlibrium and nonequilibrium conditions.
1 Introduction
Despite advances in enhanced equilibrium and path sampling methods [1, 2, 3, 4, 5, 6, 7], the study of complex biomolecular systems remains resource-intensive, often surpassing researchers’ computational capabilities. Conventional molecular dynamics (MD) cannot produce well-sampled conformational distributions except in special cases [8]. A major challenge is estimating the “steady state” or stationary distribution of the sampled states [9, 10, 11]. The accuracy of the stationary distribution is crucial because it reveals thermodynamic and kinetic properties of the system. The stationary distribution is needed to calculate free energies as well as identifying mechanistic pathways, rate constants, and committor reaction coordinates [12, 13, 14].
Recently, methods based on AlphaFold have generated protein structural ensembles [15, 16, 17, 18]. However, these methods are heuristic in the sense that they are not designed to produce equilibrium ensembles conforming to the Boltzmann factor. The algorithm described here, however, can leverage heuristic ensembles as the starting point for conventional MD and subsequent Boltzmann-factor reweighting. While heuristic starting configurations can also be updated with enhanced sampling [19], all such approaches can be improved by fixing the sample weights.
One principled way to estimate the steady-state distribution from sampled trajectory data is to build a “Markov state model” (MSM), i.e., a discrete-state transition matrix that leads to an approximation of the stationary distribution [20, 21]. However, previous work has noted that MSM estimates of stationarity are biased by the trajectory training data [22]. To ameliorate this issue, a further step can be taken in a MSM pipeline, namely, the trajectories can be reweighted based on the matrix’s stationary solution [13]. This procedure — referred to here as “single-shot reweighting” — can be helpful. However, it cannot correct the weights of trajectories within the chosen discrete states, nor can it provide a set of weights consistent with a subsequently computed transition matrix [12, 23], potentially skewing the estimation of observables.
Unbiased estimation only occurs in the traditional MSM framework when the trajectories in each discrete state are locally consistent with the stationary distribution for the chosen boundary conditions. For example, source-sink boundary conditions lead to the challenging requirement to sample from the nonequilibrium steady state [12, 13, 24]. The need for unbiased equilibrium or nonequilibrium samples creates a chicken and egg problem which can be addressed by an iterative solution [12, 23], at the cost of generating long trajectories. We note that non-traditional MSMs can be used to obtain unbiased estimates of observables [14], but traditional MSMs are biased for both equilibrium and nonequilibrium observables even when significant trajectory data is available [22, 25].
As an approach for correcting standard MSMs, we here introduce the “Randomized ITErative trajectory reWeighting” (RiteWeight) algorithm. RiteWeight reweights trajectories generated without biasing forces into their correct stationary distribution when sufficient data is available. The method applies to both equilibrium and nonequilibrium steady states. It naturally uses trajectories of any length, as short as just a single time step, as it does not rely on dynamical relaxation. Thus, RiteWeight can employ data generated from standard molecular dynamics or path sampling approaches, so long as no biasing forces have been used. We emphasize that RiteWeight is distinct from conventional importance sampling, which requires a known and well-sampled initial distribution [26, 27].
Figure 1 shows how the RiteWeight algorithm iteratively reweights trajectories using the stationary measure for a discrete-state transition matrix , which is determined in each iteration by a new random clustering. The relative weights of the trajectory segments in each cluster are fixed during a single iteration. In the equilibrium case, the transition matrix used is akin to a MSM based on weighted trajectories. In the nonequilibrium case, the matrix is computed from weighted trajectories conforming to source-sink boundary conditions, akin to a “history augmented” MSM [28, 29, 25, 24]. However, RiteWeight uses no history information and thus eliminates the requirement that all segments to be traced back to the source state. RiteWeight also does not require the Markov property to hold at the cluster level, because discrete states are used only to estimate stationarity and not to propagate dynamics over time.
RiteWeight differs from prior strategies for reweighting sampled trajectory data. For instance, one approach explicitly approximates the configuration-space density function in order to reweight configurations [30]. Another method updates the weights using a variational principle for the equilibrium distribution [31]. RiteWeight, however, does not require density estimation and only uses standard Markov models without additional fitting parameters or assumed functional forms. In prior work, our research group has applied iterative reweighting without changing cluster definitions [23], but this approach does not yield the quasi-continuous distribution which is achieved using RiteWeight.
The present report evaluates the RiteWeight algorithm through mathematical analysis and empirical testing. First, we present mathematical analysis identifying the fixed point of the RiteWeight algorithm. Then, we present numerical tests based on both synthetic molecular dynamics (SynMD) and true MD trajectories of the Trp-cage miniprotein. SynMD consists of trajectories generated from a fine-grained MSM with each state mapped to an atomistic conformation, and it enables comparison to an exactly calculable reference distribution [32]. The true MD data is a single 208 s trajectory generated by the Shaw group [8]. We compare the performance of RiteWeight, single-shot reweighting, and traditional MSMs for estimating the stationary distribution from mis-distributed data sets. Overall, RiteWeight yields better agreement with reference values for all observables considered in both equilibrium and nonequilibrium scenarios. RiteWeight achieves good performance using extremely short trajectory segments, enabling computation of mechanistic, path-based observables. Last, RiteWeight results are independent of the number of clusters used to discretize the phase space.
2 RiteWeight Algorithm
The RiteWeight algorithm estimates steady-state probability distributions using kinetic information extracted from numerous short MD trajectories or one long trajectory. The algorithm processes sequential pairs of phase-space points generated by unbiased dynamics and separated by a fixed lag time: . We refer to these pairs as “trajectory segments”. Unlike MSMs, the algorithm’s ability to estimate stationary probabilities is not constrained by the choice of lag time [21, 20]. Unlike single-shot reweighting, which requires within-cluster stationarity for unbiased results [12, 13, 24], RiteWeight imposes no such constraints. Therefore, transition pair data can be collected right from the start of the trajectories, ensuring that no data is discarded and all available information is leveraged in the analysis.
2.1 The algorithm
The following steps define the RiteWeight algorithm. In this algorithm, the user must choose the number of clusters and identify a featurization that impacts the cluster definitions.
1.
Introduce features of each configuration that satisfy rotational and translational invariance. For example, Cα pairwise distances are commonly used in conventional MSMs [33]. Define a distance between configurations based on the chosen features.
2.
Assign an initial weight to each trajectory segment connecting two consecutive configurations, using the normalization . A possible choice is uniform weights , where is the number of segments.
3.
Randomly select unique configurations as cluster centers. Define clusters by mapping each configuration to the closest center using the chosen distance.
4.
Compute a transition matrix based on the current weights of the trajectory segments and the current definitions of the clusters:
(1) where means that trajectory segment begins in cluster , and means that segment terminates in cluster .
5.
Calculate the stationary probabilities for the discrete states using the left leading eigenvector of the matrix , i.e., .
6.
Define a new weight for each trajectory segment based on the current weights and the stationary solution from Step 5. For each segment starting from cluster , the new weight is given as:
(2) where is the previous weight of the trajectory segment, is the new weight, is the current iteration’s estimate of the stationary probability for the cluster , and .
7.
Repeat steps 3-6 until a user-defined convergence criterion and average the weights over the final iterations.
Several points are noteworthy. In step 2, there are two approaches for assigning the initial weights to trajectory segments. One option is to assign uniform weights to all segments, which serves as an uninformative prior. Alternatively, if there is prior knowledge regarding the stationary distribution of the system, this information can be used to assign the initial weights in a more informed manner. Random clusters are defined in step 3, and a new clustering is introduced at every iteration. In step 6, the weight of each trajectory in cluster is updated to be the currently estimated stationary probability multiplied by a fraction , where is the previous estimate for the trajectory segment and is the previous estimate for the cluster.
Last, we define convergence in the numerical experiments by running RiteWeight for – iterations, until the weights or mean first passage time estimate exhibits small, steady fluctuations. Then, we average the weights over the final 1,000 or 10,000 iterations to further reduce noise. See supplementary Figs. 7, 8, and 10 for examples of RiteWeight convergence analysis.
2.2 Analysis of RiteWeight fixed point
The mathematical analysis of RiteWeight illustrates the strengths and limitations of the algorithm. We analyze the fixed point in detail in the Appendix, and we summarize the findings here.
For simplicity and clarity, our analysis assumes that the configurations all belong to a discrete state space of “microstates” indexed by or . We expect that RiteWeight in continuous space behaves similarly to RiteWeight in discrete space at a level of resolution automatically determined by the algorithm and its stopping criterion.
In the discrete setting, RiteWeight defines each update using coarse random clusters containing multiple microstates. However, independent of the number of clusters, the fixed point of RiteWeight is determined by the stationary measure of the microstate transition matrix with elements
Here, is the initial weight assigned to segment . The Appendix proves the stationary vector of defines the fixed point of RiteWeight assuming this stationary vector is unique and the clusters are chosen according to a random hyperplane tesselation.
The fixed point analysis suggests that RiteWeight will converge to the true stationary distribution given sufficiently dense and unbiased local sampling. In the limit of high transition counts and small microstates, the matrix approaches the true transition operator. Thus, RiteWeight will find the correct distribution regardless of the number of clusters or the lag time. However, in more realistic settings, the accuracy of RiteWeight is constrained by the data analyzed, pointing to a valuable role for the adaptive sampling methods that will be mentioned in the Discussion section.
Given the description of the RiteWeight fixed point, it is natural to ask whether the RiteWeight stationary distribution can be derived directly from a fine-grained MSM transition matrix. In practice, this is not possible because the resolution of the microstates corresponding to RiteWeight cannot be easily determined for an arbitrary data set. The Results section will show the limitations of simply using a fine-grained Markov state model without RiteWeight.
3 Test systems
Here we introduce the synthetic MD and atomistic MD systems that we used in our empirical RiteWeight experiments.
3.1 Synthetic MD for Trp-cage
The first test system for RiteWeight is synthetic MD (SynMD). SynMD offers some of the complexity of atomistic proteins, but it has computational and conceptual advantages. Using this system, long trajectories can be efficiently generated, and stationary properties can be calculated exactly as a reference [32].
Model: The SynMD model for the Trp-cage miniprotein [32] is derived from an MSM trained on a 208 s atomistic MD trajectory [8]. The MSM uses 10,500 states, which is a finer discretization than a typical MSM, and each state is mapped to a specific atomistic configuration. Kinetic Monte Carlo then yields a “synthetic” MD trajectory that closely matches the statistical behavior of the original atomistic MD [32]. The MSM lag time and hence the interval between synMD configurations is 1 ns, and the mapped configurations can be processed using standard MSMs or RiteWeight. Because the synMD model is governed by an explicitly known MSM transition matrix, we can calculate the equilibrium distribution explicitly using the leading left eigenvector of the matrix.
Trajectory preparation: We prepared a data set of 5 ns SynMD trajectories using a sampling measure far from equilibrium. First we applied tICA with a 5 ns lag time [25] to the set of minimal residue-residue distances, which are defined by the closest distances between the heavy atoms of two residues separated in sequence by at least two neighboring residues. Then we assigned indices to the 10,500 MSM states based on the slowest time-lagged independent component (TIC1) [32]. We partitioned the SynMD indices into 500 bins and selected 20 initial states randomly with replacement from each bin. For each selected state, we initiated a short trajectory of 5 ns, i.e., 5 steps. Hence the unprocessed trajectory data consists of 10,000 trajectory segments, each 5 steps long, distributed roughly evenly along the TIC1 coordinate.
RiteWeight Analysis: We applied RiteWeight in a manner blind to the discrete nature of synMD, using only the mapped atomistic configurations. We defined features using the slowest 10 tICs, and we rescaled the tICs based on commute distances. Then we performed RiteWeight using either or clusters based on the rescaled tICs.
3.2 Atomistic MD for Trp-cage
The second test system for RiteWeight is a 208 s simulation of Trp-cage in explicit solvent [8]. We used this atomistic MD trajectory data to investigate equilibrium and nonequilibrium steady states, as well as path-based observables including the mean first-passage time and net fluxes.
3.2.1 Equilibrium distribution
Trajectory preparation: Again we prepared a data set of short trajectories using a sampling measure far from equilibrium. First we applied tICA with a 10 ns lag time [25] to the set of minimal residue-residue distances. Then we projected all conformations from the 208 s trajectory onto TIC1 and grouped them into 100 uniformly spaced bins. We subsampled 10,000 conformations within each bin; however, if fewer than 10,000 conformations were available in a bin, we selected all the conformations. We used these conformations as the starting points for segments consisting of two snapshots from the long trajectory separated by a 10 ns lag time.
RiteWeight and MSM Analysis: We retained a sufficient number of tICs to explain 95% of the total variance, and we rescaled the tICs based on commute distances. For RiteWeight, we performed random clustering () using the rescaled tICs as input features. For the MSM analysis, we performed -means clustering based on – clusters. For all downstream MSM analysis following clustering, we used the PyEMMA software package [34].
3.2.2 Nonequilibrium Analysis
Trajectory preparation: We harnessed the complete equilibrium-like 208 s MD trajectory to generate trajectories for nonequilibrium steady state analysis. No subsampling of the trajectory data was performed. Instead, we extracted all possible two-step segments based on a lag time of 100, 10, 1, or 0.2 ns (the shortest lag time in the original MD data).
Macrostate definitions: Following prior work [25], we used the fuzzy spectral clustering method PCCA++ with a lag time of 100 ns to probabilistically assign configurations to folded and unfolded states. Then, we defined the “intermediate” region as the 10% of configurations whose membership scores were closest to 50% folded / 50% unfolded. We assigned the remaining 90% of configurations to either the folded or unfolded region based on the membership scores. These macrostates were used for both MSM and RiteWeight analysis.
MSM Analysis: Optimal MSM hyperparameters were previously identified for the 208 s MD data set [20, 35]. We used the PyEMMA software package [34] and replicated the optimal parameter choices, including features based on minimal residue-residue distances, dimensionality reduction using tICA with 100 tICs at a lag time of 10ns (with tICs rescaled using commute distances), and clustering using -means with .
RiteWeight Analysis: In each RiteWeight iteration, we performed clustering in the intermediate region using random clusters based on the same tICA coordinates used for MSM analysis. We then constructed a transition matrix that included the intermediate clusters along with the source and sink. To enforce the non-equilibrium steady-state (NESS) condition, we applied a boundary condition during the construction of the transition matrix. Specifically, the transition probability from the sink to all other clusters was set to zero, except for the source-state cluster which was set to one. We used the same source and sink definitions for RiteWeight and MSM analysis.
4 Results
Here we present the results of applying RiteWeight to the SynMD and atomistic MD systems introduced in Sec. 3.
4.1 Synthetic MD for Trp-cage
Fig. 2 shows that RiteWeight recovers the true equilibrium distribution for the SynMD Trp-cage system starting from trajectory data that is strongly mis-distributed. Although the algorithm employs discrete MSM-like stationary solutions at each iteration, it learns a quasi-continuous distribution.
RiteWeight’s accuracy is equally high using 10 or 1,000 clusters, underscoring the algorithm’s robustness. However, RiteWeight’s convergence speed does depend on the number of clusters, with more clusters yielding faster convergence. See the KL-based convergence analysis in supplementary Fig. 7. Following this analysis, we used iterations for RiteWeight with 1000 clusters and iterations for RiteWeight with 10 clusters. In both cases, the trajectory weights reported here are averages over the final 1,000 iterations.
In contrast to the converged RiteWeight estimates, the single-shot reweighting that occurs during the first RiteWeight iteration exhibits discrepancies due to discretization error. Even with 1,000 clusters, single-shot reweighting does not accurately predict the true stationary probabilities of the Trp-cage SynMD microstates. The discrepancy here can be attributed to the lack of local equilibrium within the clusters: local equilibrium is necessary for the success of single-shot reweighting but not for RiteWeight.
4.2 Atomistic MD for Trp-cage: Equilibrium and nonequilibrium steady states
RiteWeight also recovers the correct stationary distribution for atomistic Trp-Cage both in equilibrium and nonequilibrium settings.
Fig. 3 shows estimates of the equilibrium distribution calculated using RiteWeight and MSMs. The reference values come from a direct computation using the full 208 s Trp-Cage trajectory, and the standard errors are computed by dividing the MD trajectory into three 70 ns blocks and comparing estimates. RiteWeight matches closely with the reference, despite starting from a distribution far from equilibrium. Here, the RiteWeight distribution is an average over the final 10,000 iterations in a run of 100,000 iterations. See the KL-based convergence analysis in supplementary Fig. 8.
By contrast, MSM-based reweighting fails to recover the equilibrium distribution, even when using the same features as RiteWeight with as many as 50,000 clusters. The inability to recover the correct stationarity distribution is due to an unsatisfied assumption that weights are locally in equilibrium within each cluster. The poor performance persists even when the MSM lag time is set to 100 ns; see supplementary Fig. 9.
In the nonequilibrium setting, Fig. 4 shows that RiteWeight is similarly successful. In this case, the reference stationary measure comes from modifying the 208 s Trp-cage trajectory by removing configurations that were more recently in the folded than the unfolded state. Again RiteWeight corrects the imbalances in the starting distribution and delivers a faithful copy of the stationary distribution. Here, the RiteWeight distribution is an average over the final 10,000 iterations in a run of 400,000 iterations, based on the mean first passage time convergence analysis in supplementary Fig. 10.
4.3 Atomistic MD for Trp-cage: Mean first-passage time
Next we used RiteWeight to calculate the mean first-passage time (MFPT) for atomistic trp-Cage folding. We ran RiteWeight using source-sink boundary conditions with the unfolded region as the source and the folded region as the sink. We then evaluated the MFPT using the reciprocal flux relation [36, 37]:
where is the lag time for the trajectory segments and is the estimated probability of each trajectory segment . For comparison, we also performed an MSM analysis using the first-step relation:
which is implemented in the PyEMMA software [34].
Fig. 5 shows that RiteWeight recovers the reference MFPT values, even with a short lag time of ns. We note that the reference MFPT from the 208 s trajectory slightly increases with lag time, because some of the first entries to the target folded state are missed at longer lag times. Consistent with previous work [25], the MSM recovers the correct MFPT but only at sufficiently long lag times ns. At short lag times ns, the MSM underestimates the MFPT by an order of magnitude or more.
RiteWeight is able to reproduce the MFPT because it uses a self-consistent iteration that locally corrects the weights within each cluster. In contrast, an MSM implicitly uses uniform weights within each cluster without correction. The initial trajectories are distributed close to equilibrium, and the MSM does not account for the nonequilibrium distribution from the source-sink boundary conditions that is “tilted” with respect to equilibrium [38]. We conjecture that updating the weights is especially important at short lag times, since the trajectories contributing to the MFPT estimate are the peripheries of the clusters which is the region of greatest tilting.
4.4 Atomistic MD for Trp-cage: Net fluxes
A major goal of molecular simulation is understanding the “mechanisms” of a functional transition, here defined as the temporal sequences of molecules. We analyzed the mechanisms for atomistic Trp-cage folding using net fluxes:
where the cluster indices and are ordered so that the net flux is positive in the reference MD data. For comparison, we also performed an MSM analysis using the formula for the net fluxes involving the backward committor [25]:
Here is the equilibrium steady state, which comes from solving , and is the backward committor vector computed by pyEmma [34], which gives the probability that a Markov chain with transition matrix was more recently in the unfolded state than the folded state.
Fig. 6 shows that the net fluxes from RiteWeight closely match the MD reference for each lag time 0.2, 1, 10, or 100 ns. Thus, RiteWeight precisely describes the transition processes leading to folding, even processes with a lag time ns which is the shortest lag time available in the original MD data. In contrast, the net fluxes from MSMs exhibit large discrepancies, echoing a previously identified deficiency of the MSM approach [25]. At each lag time, the MSMs exhibit some net fluxes in the opposite direction of MD, implying a different temporal sequences of events. The MSM fluxes are most quantitatively accurate at the 100 ns lag time, which is 1-2 orders of magnitude longer than the transition-path time over which folding events typically take place [25]. Supplementary Figs. 11 and 12 report confidence intervals for the MD reference fluxes, revealing tht a range of MSM estimates fall outside the 95% confidence intervals at each lag time 10 ns.
5 Discussion and Conclusions
We have introduced RiteWeight, an algorithm for correcting distributions that deviate from equilibrium or nonequilibrium steady states. Although it is built on a framework of Markov state models (MSMs) that discretize phase space, RiteWeight iteratively achieves self-consistency among all possible discretizations. This new approach overcomes a key limitation of MSMs by correcting the distribution within each discrete region to conform to the steady state. It accurately evaluates the steady state without requiring the Markov property at the discrete level. When the internal distributions match the steady state, a wider array of observables — such as mean first passage times and net fluxes — can be calculated with high accuracy even at short lag times [25]. The robustness of RiteWeight is illustrated by the insensitivity to the number of clusters used during the iteration process (Fig. 2).
RiteWeight produces models analogous to “history augmented” MSMs [39, 38, 28, 25] but with a key difference: no history information is used to analyze the data. For this reason, RiteWeight is capable of processing a wide variety of dynamics data, even data from adaptive sampling [40, 41]. The raw data can be generated from multiple trajectories of arbitrary lengths initiated from arbitrary starting points.
Building on previous reports [25], we have probed the shortcomings of MSMs for identifying mechanisms of folding in the atomistic Trp-cage system (Fig. 6). Surprisingly, the MSMs built from MD reference data continue to show discrepancies from reference values even at lag times of 100 ns. At sufficiently long lag times, the system must become fully Markovian, but evidently at 100 ns, the net fluxes are still influenced by the initial distributions within the clusters. This sensitivity of the net fluxes contrasts with the relative robustness of MSM mean first passage time estimates that match the reference value at a lag time of 100 ns (Fig. 5).
There remain important tasks ahead to realize the full potential of RiteWeight. Most notably, the data analyzed here consists of large data sets with dense sampling in visited regions of phase space. In sparser data sets, the algorithm may need to rely on other regularization and smoothing strategies. For example, we can modify the RiteWeight update formula (2) to read
Here, is a learning rate hyperparameter within the interval . Choosing a smaller value may mitigate the impact of noise without changing the RiteWeight fixed point (Theorem 1 in the Appendix). Another valuable strategy that may improve the estimation of observables in key phase-space regions is adaptive sampling [40, 41]. Finally, because RiteWeight can be accurate at any lag time, the range of lag times could be explored to optimize numerical efficiency in equilibrium calculations.
We foresee diverse opportunities for applying RiteWeight in the future. To promote the longstanding goal of generating Boltzmann-weighted ensembles for proteins, RiteWeight could be applied to MD trajectories initiated with machine-learning approaches based on AlphaFold [15, 16, 17, 18] or trajectories initiated with nuclear magnetic resonance data [42, 43]. Further, RiteWeight could be used to iteratively calculate local mean first passage times and generate transition paths in combination with the weighted ensemble method [44, 45].
In conclusion, the new RiteWeight algorithm solves the problem of reweighting flawed initial data sets into desired steady state distributions. The iterative use of random clusters leads to quasi-continuous and accurate final distributions. Applications to equilibrium and nonequilibrium observables compare favorably to state-of-the-art methods, with a particular advantage in quantifying nonequilibrium net fluxes.
Data Availability Statement
The python package for running the RiteWeight algorithm is available at: https://github.com/ZuckermanLab/rite_weight. The data used in this study to evaluate RiteWeight’s performance is available upon request. The SynMD Trp-cage model used in this study is hosted at: https://github.com/jdrusso/SynD-Examples
Acknowledgments
We thank Jeremy Copperman and John Russo for valuable discussions on the subject of reweighting trajectories. The authors are grateful for financial support from the NIH under Grant GM115805. We thank DE Shaw Research for the Trp-cage trajectory data.
Appendix: Analysis of RiteWeight fixed point
This section identifies the “fixed point” of the RiteWeight algorithm, which is defined as follows.
Definition 1 (Fixed point).
A collection of trajectory weights with is a fixed point of RiteWeight if the weights stay the same during each RiteWeight iteration for any positive probability choice of partition.
In practice, we typically apply RiteWeight using a continuous state space and clusters defined by a random Voronoi tesselation. Here for simplicity we consider a finite state space with clusters defined by a random hyperplane tesselation. See [46] for background on random hyperplane tesselations, which are called “stable under iteration tesselations” in stochastic geometry.
Assumption 1 (Finite state space).
The state space consists of a finite number of distinct microstates .
Assumption 2 (Random hyperplane tesselation).
The RiteWeight clusters are generated through one or more iterations of the following procedure. Initially, there is a single cluster containing all the microstates. At each iteration, any cluster that contains at least two microstates is randomly split by a hyperplane into two new clusters as follows. First, we choose a uniformly random direction that is normal to the hyperplane
Then, conditional on , we choose a uniformly random offset
The normal direction and the offset define two new clusters of microstates,
that are split by the hyperplane.
Under Assumptions 1 and 2, the following main result shows that the RiteWeight fixed point is completely determined by the microstate transition matrix.
Theorem 1 (RiteWeight fixed point).
Consider a finite state space (Assumption 1) and a collection of trajectories with weights satisfying . Define the associated microstate transition matrix with entries
and assume has a unique stationary measure. Here, means that segment begins in microstate , and means that segment ends in microstate . Then, under the random hyperplane model (Assumption 2), is a fixed point of RiteWeight if and only if the vector with entries
is the fixed point of , that is,
Proof.
First we check that equation (11) implies is a fixed point of RiteWeight. By assumption, the microstate transition matrix has a unique stationary measure, so it is irreducible and aperiodic. It follows that the cluster transition matrix is irreducible and aperiodic, so has a unique stationary measure also. Next, equation (11) implies
This confirms that is a fixed point of RiteWeight.
Next, we assume that is a fixed point of RiteWeight and check that equation (11) holds. Since is a fixed point, we must have for each cluster and consequently
By considering all positive probability random clusters , we arrive at the identity
where is the matrix whose columns are characteristic functions for each positive probability random cluster: values indicate membership in the cluster and values indicate non-membership. To complete the proof, we will show that has full column rank and therefore (11) holds.
We observe that the first iteration of the random hyperplane tesselation generates a uniformly random direction that leads to distinct values for every distinct microstate with probability one. Hence, there must be a consistent ordering of the microstates, say , so that the event
occurs with positive probability. It follows that a division into clusters
occurs with positive probability for each . With the microstates ordered in this way, the linear span of the columns of includes all the vectors , where is a standard basis vector. This shows has full column rank, completing the proof. ∎
Supplementary figures